if there are currently 45 heavy lifters at factory a and company increases it's fleet by 60% , how many total lifters will be at factory a and factory b after they are redistributed


1
Expert's answer
2022-01-14T09:43:40-0500
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 2009 & 2010 \\ \hline
 Factory\ A\ 15\% & Factory\ A\ 20\% \\
 \hdashline
 Factory\ B\ 15\% & Factory\ B\ 35\% \\
 \hdashline
\end{array}"

As in 2009 the Factory A had 45 lifters and it is the 15% of the total quantity therefore the total quantity was


"45\cdot\dfrac{100\%}{15\%}=300\text{ lifters}"

In 2010 the company increase it’s fleet by 60% and now it has 


"300\cdot\dfrac{160\%}{100\%}=480\text{ lifters}"

The factories A and B have the "20\% + 35\% = 55\%" of the total quantity of lifters and it’s "480\cdot\dfrac{55\%}{100\%}=264\text{ lifters}."

"264\text{ lifters}."